answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How can you choose a random sample of teachers?

The first step is to establish a sampling frame. This is a list of all teachers in the domain that you are interested in.

Next you allocate a different number to each teacher.

Then you use a random number generator to generate random numbers. You select each teacher whose number is generated. If the teacher has already been selected for inclusion in the sample, you ignore the duplicate and continue until you have a sample of the required size.

What is the probability of genotypes and phenotypes?

What is the probability that any of the offspring between individuals with the genotype AABbCcddEE will have the genotype AABBCCddEE

What is the probability of an adverse event ocurring?

There terms frequent, likely, occasional, seldom, and unlikely, used in the risk assessment matrix is the level of

What approach to probability is based on a persons degree of belief and hunches that a particular event will happen?

Subjective

If you assume particular events will happen with a certain prior distribution, that is Bayesian probability.

What is the probability of pulling two red cards without replacement from a deck of cards?

Assuming a pack consists of 52 cards as per normal.

Initially half the cards are red.

Probability that the first card drawn is red = 1/2.

Now there are 25 red cards left out of 51 remaining cards.

Probability that the second card drawn is red = 25/51.

Probability that both cards drawn are red therfore = 1/2 * 25/51 = 25/102

What is the benefit of transforming standard normal distributions to conform to the standard distribution?

There are no benefits in doing something that cannot be done. The standard normal distribution is not transformed to the standard distribution because the latter does not exist.

What is the probability that it is an ace or a heart in a 52-deck of cards?

If you add 1 for the ace to 13 for Hearts cards you have 14 cards out of a 52 card pack. So the probability of selecting one of those 14 cards is 14/52 = 7/26

That is 0.2692307692307.........( recurring 69230)

Correcting to 4 significant figures that equals 0.2692 which equals 26.92%

That means that you will select a correct card in just over 1 in 4 times.

Most winning lottery numbers?

There is insufficient information for us to answer this question. Please edit the question to include more context or relevant information. It would be useful to know which lottery from which country the question refers to.

If you draw a card from a standard deck of cards what is the probability of drawing a red card or a face card?

The probability of drawing a red card is 1/2 (26/52).

The probability of drawing a face card is 3/13 (12/52).

The probability of drawing a red face card is 3/26 (6/52).

The probability of drawing a red card or a face card is 8/13 (32/52).
2 out of 52

or 1 out of 26.

What is the probability of getting 3 girls in a family of 10 children?

The probability of exactly 3 girls in a family of 10 children, assuming equal chance of a boy or girl, is 0.1172. This is a binomial distribution.

What is the most probable ion formed from Li?

The only probable ion to form from Lithium is the Li+ ion.

What was the major event that occurred in 2000?

The disputed election for President of US, between Albert Gore and George W Bush.

Which Queen is depicted on playing cards?

The queen on a standard modern deck of Anglo-American playing cards depicts no one in particular, and is most likely a distorted version of older versions of European cards.

Older sets of cards typically varied based on the origin of the card. Face cards typically depicted historical figures of religious or political importance.

Source: Wikipedia

How do you identify a normal curve?

You cannot. It has a characteristic bell-shaped curve but so does a Student's t with enough degrees of freedom. There are other distributions which, with suitable choice of parameters can be made to look very similar to the Normal curve.

When rolling 1 dice what is the probability of rolling number 4?

It is 1 (a certainty) if you roll it often enough. For a single roll of a fair die, the probability is 1/6.

What is probability of a heart or a club?

In a standard deck of 52 cards - the probability of drawing any single card of two suits is 1:2 or 50%.

What is the chance of getting a password right with random letters?

Around 65-70%. I am a expert on the subject and have preformed many experiments where I have someone type random letters and see the outcome. Most of the time it is right somehow.

Trending Questions
Where can one get a rolling chassis? Scores on an exam are normally distributed with a mean of 76 and a standard deviation of 10 In a group of 230 tests how many students score below 66? If two fair dice are rolled what is the probability that the total number of spots shown is divisible by 2 or 3? What does DX Error 0x8876086c mean? How many meaningful English words can be formed with the letters TPO using each letter only once in each word? How many cubic yards are in a sidewalk that's 48in wide by 40ft long by 3.5 in thick? If the sample mean is 10 the hypothesized population mean is 9 and the population standard deviation is 4 compute the test value needed for the z test? What probability is based on knowing all of the equally likely outcomes of an experiment? What probability of rolling a 6 on the eighth toss of a die? Is tossing 1 coin 100 times the same as tossing 100 coins 1 time? A single die is toss Find the probability of getting a 6? If you draw 5 balls with replacement from an urn containing white balls and black balls is the probability of getting 3 white balls always higher than if you draw the balls without replacement? How many twos are there in a standard deck of cards? What were the outcomes of mau mau? If you have two dice how many products appear? Paolo has 40 marbles in a bag. If he splits up the marbles among his 4 friends he'll have 3 left over? Is probability independent? When is the next geomagnetic reversals? You draw a card replace it then draw another card.? How many possible outcomes exist when rolling a number cube that has sides that say 3 5 and 6?